OpenStudy (anonymous):

Solve the following equation: 8x^2=3-12x

OpenStudy (luigi0210):

move and set equal to zero

OpenStudy (anonymous):

I got: -3+- square root 15/4

OpenStudy (anonymous):

8x^2+12x-3=0 Now use quadratic formula to get the roots. I hate using the equation options on here, so I'll just let you handle that =P

OpenStudy (anonymous):

And also, your answer sounds about right.

OpenStudy (anonymous):

i. hate. quadratic. formula.

OpenStudy (anonymous):

I do too, it's like forcing myself to eat broccoli and asparagus... cereal

OpenStudy (luigi0210):

Yup that's right: \[-\frac{ 3+/-\sqrt{15} }{ 4 }\]

OpenStudy (anonymous):

sweet magic.:)